OpenStudy (anonymous):

I need help with volume in geometry. Please help? When you put an object into a container of water, it raises the water level 3 cm. The container is a rectangular prism whose base measures 10 cm by 12 cm. Without the object in the container, the water is 8 cm high. a. What is the volume of the container with just the water in it? b.What is the volume of the container with the water and object in it? c.What is the volume of the object?

OpenStudy (jackofalltradez):

a.) 10 * 12 * 8 = 960 cm cubed b.) 10 * 12 * 11 = 1320 cm cubed c.) 1320 - 960 = 360 cm cubed

OpenStudy (jackofalltradez):

Hope that helped, @Nyan64 :)